OpenStudy (anonymous):

Evaluate each polynomial for x = 3 and for x = −3. −x^2 − 3x^3 − x^4

OpenStudy (anonymous):

Plug in x to be 3 and see what you get.. then plug in -3 and see what you get.

OpenStudy (anonymous):

okay so far this is what I got (-3*-3) - 3(-3*-3*-3) - (-3*-3*-3*-3) 9 - -81 - 81 = 153

OpenStudy (anonymous):

that was for -3

OpenStudy (anonymous):

For -3, you should get -9

OpenStudy (anonymous):

3^2 - 3(3^3) - 3^4 is that right so far

OpenStudy (anonymous):

Well maybe i did something wrong.. Just a sec.

OpenStudy (anonymous):

3*3 - 3*3*3*3*3 - 3*3*3*3= -315

OpenStudy (anonymous):

No,, for x=-3 you should get -9

OpenStudy (anonymous):

okay got that one

OpenStudy (anonymous):

You have.. - 3*3 - 3*3*3*3 - 3*3*3*3

OpenStudy (anonymous):

the one I just posted is for 3

OpenStudy (anonymous):

For x = 3 you should get -171

OpenStudy (anonymous):

okay so if I plug in -3*-3 then it should come out to -171

OpenStudy (anonymous):

If you plug in 3 you get -171

OpenStudy (anonymous):

it its -3 then you get -9

OpenStudy (anonymous):

how come then when I do it on the calculator it comes out to be -315

OpenStudy (anonymous):

Your not doing the parenthesis right.. type it in like this\[(3)^2 - 3(3)^3 - (3)^4\]

OpenStudy (anonymous):

okay tell me if I am doing this right 3*3=9 3*3*3=27*3=81 3*3*3*3=81 right

OpenStudy (anonymous):

9-81-81 = 153

OpenStudy (anonymous):

-(3*3) - 3*3*3*3 - 3*3*3*#

OpenStudy (anonymous):

okay forgot the -

OpenStudy (anonymous):

-9 - 81 - 81

OpenStudy (anonymous):

okay now I got it thank you